\begin{thebibliography}{00}

% 6
\bibitem{Balick74a} B. Balick and R. L. Brown 1974 Intense sub-arcsecond structure in the galactic center \textit{Astrophys. J.} \textbf{194} 265--270

% 9
\bibitem{Brown82a} R. L. Brown 1982 Precessing jets in Sagittarius A: gas dynamics in the central parsec of the galaxy \textit{Astrophys. J.} \textbf{262} 110--119

% 15
\bibitem{Abbott16a} B. P. Abbott \textit{et al.} 2016 Observation of gravitational waves from a binary black hole merger \textit{Phys. Rev. Lett.} \textbf{116} 061102

% 16
\bibitem{Issaoun19a} S. Issaoun \textit{et al.} The size, shape and scattering of Sagittarius A* at 86 GHz: first VLBI with ALMA \textit{Astrophys. J.} \textbf{871} 30

% 4
\bibitem{Williams71a} E. R. Williams, J. E. Faller and H. A. Hill 1971 New experimental test of Coulomb's law: a laboratory upper limit on the photon rest mass \textit{Phys. Rev. Lett.} \textbf{26} 721--724

% 8
\bibitem{Chibisov76a} G. V. Chibisov 1976 Astrophysical upper limits on the photon rest mass \textit{Sov. Phys. Usp.} \textbf{19} 624--626

% 13
\bibitem{Adelberger07a} E. Adelberger and G. Dvali and A. Gruzinov 2007 Photon-mass bound destroyed by vortices \textit{Phys. Rev. Lett.} \textbf{98} 010402

% 1
\bibitem{Proca36a} A. Proca 1936 Sur la th\'{e}orie ondulatoire des \'{e}lectrons positifs et n\'{e}gatifs \textit{J. Phys. Radium} \textbf{7} 347--353 

% 3
\bibitem{Goldhaber71a} A. S. Goldhaber and M. M. Nieto 1971 Terrestrial and extraterrestrial limits on the photon mass \textit{Rev. Mod. Phys.} \textbf{43} 277--296

% 11
\bibitem{Jackson01a} J. D. Jackson 2001 \textit{Classical Electrodynamics} (John Wiley \& Sons)

% 2
\bibitem{Stueckelberg38a} E. C. G. Stueckelberg 1938 Die Wechselwirkungs Kraefte in der Elektrodynamik und in der Feldtheorie der Kernkraefte \textit{Helv. Phys. Acta} \textbf{11} 225-244, 299--312, 312--328

% 5
\bibitem{Bekenstein72a} J. D. Bekenstein 1972 Nonexistence of Baryon number for static black holes \textit{Phys. Rev. D} \textbf{5} 1239--1246

% 14
\bibitem{Dolgov07a} A. Dolgov and D. N. Pelliccia 2007 Photon mass and electrogenesis \textit{Phys. Lett. B} \textbf{650} 97--102

% 7
\bibitem{Vilenkin78a} A. Vilenkin 1978 Self-interaction of charged particles in the gravitational field \textit{Phys. Rev. D} \textbf{20} 373--376

% 10
\bibitem{Leaute85a} B. Leaute and B. Linet 1985 Self-interaction and massive vector field in the Schwarzschild space-time \textit{Gen. Rel. Gravit.} \textbf{17} 783--798

% 12
\bibitem{Paul04a} A. Paul 2004 Time scale for loss of massive vector hair by a black hole and its consequences for proton decay \textbf{70} \textit{Phys. Rev. D} 124005

\end{thebibliography}

\documentclass[prd,twocolumn,amsmath,amssymb]{revtex4-1}
\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{MnSymbol}
\usepackage{mathrsfs}
\usepackage{colortbl}
\makeatletter

\begin{document}

\title{Using a black hole to weigh light: \\ can the Event Horizon Telescope yield new information about the photon rest mass?}

\author{Robert P. Cameron}
\email{robert.p.cameron@strath.ac.uk www.ytilarihc.com}
\address{SUPA and Department of Physics, University of Strathclyde, Glasgow G4 0NG, U.K.}

\begin{abstract}
\noindent We point out that data collected by the Event Horizon Telescope or a similar project might yield new information about the photon rest mass $m_\gamma$, in the form of evidence that $m_\gamma\ne 0$ together with a lower bound on $m_\gamma$ or a new upper bound on $m_\gamma$. Using Sgr A*, there is scope to improve on the best upper bound obtained via laboratory tests of Coulomb's law by up to nineteen orders of magnitude.
\end{abstract}

\date{\today}
\maketitle


It is an exciting time for the study of black holes. The Event Horizon Telescope is due to release the first image of an event horizon, namely that of the black hole Sgr A* at the centre of our galaxy \cite{Balick74a, Brown82a}. This comes hot on the heels of the first direct detection of a binary black hole merger, using gravitational waves \cite{Abbott16a}. 

In this paper we point out that data collected by the Event Horizon Telescope or a similar project \cite{Issaoun19a} might yield new information about the photon rest mass $m_\gamma$.

Surprisingly, it is not yet known with certainty how much light weighs: it is usually assumed that $m_\gamma$ is \textit{exactly} zero, however this has never actually been confirmed. The most precise laboratory tests of Coulomb's law have only succeded in establishing that $0\le m_\gamma\lesssim10^{-50}\textrm{kg}$ \cite{Williams71a} and the strongest empirical claim made to date, on the basis of astronomical observations, is that $0\le m_\gamma\lesssim10^{-63}\textrm{kg}$ \cite{Chibisov76a} (the reliability of the assumptions underlying this claim has been disputed \cite{Adelberger07a}). The possibility remains that $m_\gamma \ne 0$.

The question of whether or not $m_\gamma$ is indeed exactly zero is of fundamental importance. If it were discovered that $m_\gamma$ has a non-zero value (no matter how small), Maxwell's equations would have to be replaced in principle by the Proca equations \cite{Proca36a, Goldhaber71a, Jackson01a}. This would dramatically alter our basic understanding of light: according to the Proca equations, the potential $A^\mu$ is uniquely defined and thus directly observable; light has three possible polarisation states, not two; light does not propagate at the universal speed limit $c$ but instead with phase and group speeds that differ from each other and vary with frequency, even in vacuum \cite{Goldhaber71a, Jackson01a}. A non-zero value for $m_\gamma$ would also imply the existence of a new elementary scalar field or fields to produce the mass \cite{Stueckelberg38a}.

A remarkable prediction was made by Bekenstein in 1971: a static black hole with electric charge has a Coulombic electric field outside the event horizon if $m_\gamma=0$ but \textit{no} electric field outside the event horizon if $m_\gamma\ne0$, regardless of how small $m_\gamma$ is; a corollary of the no-hair theorem \cite{Bekenstein72a, Dolgov07a}. Further static calculations revealed that the electric charge of matter outside the event horizon is, in effect, \textit{screened} by the hole if $m_\gamma\ne0$, with no such screening if $m_\gamma =0$ \cite{Vilenkin78a, Leaute85a, Dolgov07a}. Finally, dynamic calculations revealed that this screening is realised after a time $\tau \sim 1/\mu_\gamma c=\hbar/m_\gamma c^2$ \cite{Paul04a, Dolgov07a}, where $\mu_\gamma=m_\gamma c/\hbar$ is the Compton wavenumber of the photon. Thus, electromagnetic radiation produced by matter just outside the event horizon of a black hole should differ dramatically depending on whether $m_\gamma =0$ or $m_\gamma \ne 0$, assuming that this matter has been in the vicinity of the hole long enough for the screening to be realised if $m_\gamma \ne 0$. The possibility of exploiting this phenomenon to extract information about $m_\gamma$ does not appear to have been pointed out explicitly before, perhaps because of the extreme difficulty of the necessary observations.

We propose that data collected by the Event Horizon Telescope or a similar project \cite{Issaoun19a} be analysed with the above in mind. There are two distinct possibilities. If the electromagnetic radiation produced by matter just outside the event horizon of a black hole reveals that screening \textit{does} occur, it can be concluded that $m_\gamma \ne 0$ and, furthermore, that $m_\gamma>\hbar/\tau c^2$, where $\tau$ is the duration for which the matter has been in the vicinity of the hole. As indicated above, the discovery that $m_\gamma\ne0$ would be revolutionary. If, instead, the electromagnetic radiation produced by matter just outside the event horizon does \textit{not} reveal any sign of screening, it can be concluded that $0\le m_\gamma \lesssim\hbar/\tau c^2$. For Sgr A*, $\tau$ could conceivably be as large as the age of the Milky Way, in which case $\hbar/\tau c^2=10^{-69}\textrm{kg}$. Thus, there is scope to improve on the best upper bound on $m_\gamma$ obtained via laboratory tests of Coulomb's law ($10^{-50}\textrm{kg}$ \cite{Williams71a}) by up to nineteen orders of magnitude and even the strictest empirical upper bound on $m_\gamma$ yet claimed ($10^{-63}\textrm{kg}$ \cite{Chibisov76a}; disputed in \cite{Adelberger07a}) might be beaten, by up to six orders of magnitude.

This work was supported by The Leverhulme Trust (RPG-2017-048).



\end{document}